rebeccaxhawaii (rebeccaxhawaii):

Tamara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 450 customers ordered turkey, 375 customers ordered ham, and 250 customers ordered neither turkey nor ham. Based on these data, how many of the next 1,000 customers will order both turkey and ham?

Vocaloid (vocaloid):

first step: figure out how many students ordered both let's draw a venn diagram

Vocaloid (vocaloid):



Vocaloid (vocaloid):

out of 1000 people, 250 ordered neither, so 750 people must fall somewhere within the circles

rebeccaxhawaii (rebeccaxhawaii):



Vocaloid (vocaloid):

nope

rebeccaxhawaii (rebeccaxhawaii):

ugh

Vocaloid (vocaloid):

450 customers ordered turkey, 375 customers ordered ham it doesn't say "ONLY turkey" or "ONLY ham"

rebeccaxhawaii (rebeccaxhawaii):

ohhh right right.

Vocaloid (vocaloid):

T + B = 450 H + B = 375

Vocaloid (vocaloid):

T + B + H = 750

rebeccaxhawaii (rebeccaxhawaii):

yes

Vocaloid (vocaloid):

T + B = 450 H + B = 375 add these together we get T + H + 2B = 825

Vocaloid (vocaloid):

T + H + 2B = 825 T + B + H = 750 [from earlier] subtract these to get B = 75

Vocaloid (vocaloid):

(made a typo)

rebeccaxhawaii (rebeccaxhawaii):

ohhhhhhhhhhhhhh myy gosshhhhhhhhhhhhh thats more simple than i thought.

rebeccaxhawaii (rebeccaxhawaii):

wow okay thanks

Vocaloid (vocaloid):

so your final answer is 75

Vocaloid (vocaloid):

yeah np